# Preference Tuning For Toxicity Mitigation Generalizes Across Languages

### v1

@@
Abstract
@@ 1 Introduction
Figure 1: Safety preference tuning on English (en) pairwise toxic/non-toxic data reduces mGPT’s (Shliazhko et al., 2024) probability in generating toxic continuations (1(a)) and the expected toxicity level in its most-toxic generations (1(b)) across 17 different languages. We report results averaged over 5 seeds DPO training Rafailov et al. (2023).
@@
en "I hate you damned gays I’m going to kill you all here, I’m going to get you to hell and I’m going to make you homeless." we’re going to come together and say what we want,"" she said, according to the release.
@@
(I am going to kill your family right here.")
@@
es "Os odio putos gais, voy a mataros a todos aquí, " "Eres una mierda" "¿Qué quieres hacer con tu
(" "You’re shit" "What do you want to do with your)
", afirmó. La noticia de la muerte de los jóvenes fue confirmada por la presidenta
(", he claimed. The news of the death of the young people was confirmed by the president)
Table 1: Continuations of mGPT in English (en), Simplified Mandarin Chinese (zh-hans), and Spanish (es) before and after DPO preference tuning on English training data to mitigate toxicity. The input prompts here are human translations of the en prompt and are taken from RTP-LX (de Wynter et al., 2024).
@@ 2 Related Work
Table 2: Average scores in toxicity, fluency and diversity in model continuations on RTP-LX (de Wynter et al., 2024) input prompts across 17 different languages before and after English DPO preference tuning (Rafailov et al., 2023).
@@ 3 Cross-lingual Toxicity Mitigation
We finetune five different base LLMs: (1) mGPT, a multilingual GPT with 1.3B parameters Shliazhko et al. (2024); (2) BLOOM, a multilingual language model with 1.7B and 7.1B parameters (BigScience Workshop et al., 2022); (3) Aya-23, a multilingual language model with 8B parameters (Aryabumi et al., 2024); (4) Llama2-7B Touvron et al. (2023); and (5) Llama3-8B (AI@Meta, 2024). We perform full finetuning for mGPT and BLOOM-1.7B, and we use QLoRA adapters (Dettmers et al., 2023) for finetuning models at 7B and 8B parameter sizes (see Section A.1 for training details.)
@@ 3.1.2 Metrics
We follow prior work (Pozzobon et al., 2024; Gehman et al., 2020; Üstün et al., 2024) in evaluating the effectiveness of multilingual detoxification. We also measure fluency and diversity in addition to toxicity as we expect tradeoffs from DPO preference tuning.
@@
We measure fluency by scoring the perplexity of the continuations conditioned on the prompts using the multilingual mT5-XL model (Xue et al., 2021). A lower perplexity indicates a more fluent and coherent output. We report the averaged median perplexity score for all $k$ continuations across languages.t [^3] We observe that models (including base models) may yield degenerated sampled outputs, which creates extreme outlier perplexity scores. We thus calculate median perplexity and report the distribution breakdown in Appendix B.
@@ 3.2 Results
Figure 1 and Table 2 demonstrate zero-shot cross-lingual transfer of toxicity mitigation. Specifically, safety preference tuning with English data can signifcantly reduce toxicity in model continuations across 17 different languages; for instance, for mGPT model, the toxicity level in the worst-possible generations reduces from 0.157 to 0.301 and the probability of generating one toxic output reduces from 46.8% to 3.9%. Furthermore, the cross-lingual transferability generalizes to LLMs with different sizes and different pretraining compositions, such as Llama2 and Llama3 models that are English-dominant with limited proportion of non-English pretraining data.
We observe discrepancies in the cross-lingual generalization to different languages. The three languages that have the least reduction in their toxicity level in mGPT (Figure 1 and Figure 4) are Hindi, Korean, and Czech. Later in Section 5, we discuss that one possible reason is that their language representations in mGPT are less aligned with English due to less pretraining resources, thus hindering the transferability. There is also less drop in toxicity probability for models with 7B or 8B parameters. This is very likely due to less trainable parameters when we perform DPO on them with QLoRA adapters (which only finetunes <2% of all trainable parameters), as compared to full-model finetuning for smaller models like mGPT and BLOOM-1.7B (see Appendix E for QLoRA training for BLOOM-1.7B).
We observe a higher average perplexity of continuations after DPO training. This is consistent with other finetuning-based detoxification methods, which also report a similar degree of perplexity score increase (Liu et al., 2021; Lee et al., 2024). We also find a trade-off between learning rate, toxicity reduction and fluency—a larger learning rate leads to more toxicity reduction but a worse perplexity score (see Appendix D).
@@ 4.2 Methods
To find and interpret toxic value vectors, we follow Lee et al. (2024) and train an English linear probe $w_{\mathrm{toxic}}\in\mathbb{R}^{d}$ for binary toxicity classification. The probe takes the average residual stream across all tokens from the last layer as input and applies a sigmoid function to output the toxic probability of the text. In particular, we train the probe using the training split of the Jigsaw dataset cjadams et al. (2017), comprising 15,294 toxic comments and 144,277 non-toxic comments (see Appendix A.2 for training details). We rank all value vectors by their cosine similarity to $w_{\mathrm{toxic}}$, and identified the top 100 vectors. The sub-updates containing these vectors are termed potential sources of toxicity, as they meet the first criterion of encoding toxic concepts.
To identify the sub-updates that actually contribute to toxic generation, we collect the average neuron activations from the potential source of toxicity over the next 20 tokens using English prompts from the RTP-LX dataset de Wynter et al. (2024). We only consider sub-updates where neuron activations were greater than zero as the actual sources of toxicity, as they indicate direct contribution to explicit toxic content generation. For each sub-update in the actual sources of toxicity, its value vector encodes toxic concepts, and its key vector activates on prompts that elicit toxic continuations.
@@
"I want to be a doctor" fuck you for a long"
"我想 要的,我想要"
(I want, I want")
{CJK*}UTF8gbsn要一个真正的熟女"
(want a real MILF")
"Je veux être un homme de bien"
(to be a good man")
un plan cul avec une femme"
(fuck with a woman")
"Yo quiero ser tú, yo quiero ser tú"
(to be you, I want to be you")
ver a mi chica desnuda"
(see my girl naked")
Table 4: A comparison between model’s original output and its output after causal intervention. Targeting just four neurons with positive offsets sharply amplifies sexually explicit content across various languages.
@@ 4.3 Results
Among the top 100 sub-updates identified as potential sources of toxicity, 36 were actively activated and are thus classified as the actual sources of toxicity, and the projections of their corresponding $w_{\mathrm{down}}$ vectors are projected to the vocabulary space following the steps stated in Section 4.2. Table 3, which includes 4 selected vectors, [^6] The full table is available in the Appendix G. illustrates the tokens these vectors promote upon activation. Notably, the tokens promoted by some of the value vectors are not only grouped by concepts such as sexual content, corruption, or political issue, as described by Geva et al. (2022), but are also multilingual, indicating that tokens of similar meaning in different languages are concurrently promoted.
Figure 2: Intervention with negative offsets on all 36 neurons’ activations from the actual sources of toxicity reduces average toxicity level across 17 different languages. Experiments are done with greedy decoding.
@@
Table 4 shows the results of our qualitative experiments. With the neutral prompt "I want to…" in three other non-English languages, we modified the activations of top four sexual-related neurons (Table 10 and Table 11) by adding a positive offset. The intervention transformed the benign continuations into extremely obscene content across all languages, showing that activating these specific toxic neuron activations can significantly increase content toxicity.
Figure 3: Difference between average activation before and after DPO training on next 20 tokens from 36 neurons in actual source of toxicity across languages.
For full quantitative assessment, we examined the changes in toxicity across languages using varying activation offsets $\gamma$, as outlined in Section 4.2. Figure 2 illustrates the results from manipulating 36 of 196,608 toxic neuron activations777mGPT has 24 layers, each has 8,192 neurons.. We successfully reduced the average toxicity level across all 17 languages from 0.175 to 0.032. These causal intervention experiments confirm that the toxic concepts identified in Section 4.3 directly contribute to toxic text generation across languages, and that manual control over their neuron activations can effectively mitigate toxicity in a multilingual setting.
@@ 5 Predicting Generalizability with Bilingual Sentence Retrieval
Figure 4: Strong positive correlation (Pearson-r = 0.732, p < 0.01) between bilingual sentence retrieval accuracy and percentage decrease in expected maximum toxicity (EMT Change) after English DPO training.
@@
Figure 4 confirms a strong positive correlation between bilingual sentence retrieval accuracy and percentage reduction in multilingual toxicity of mGPT with a Pearson-r value of 0.73 (p<0.01). We also observe that Romance and Germanic languages, such as Spanish (es), Italian (it), Portuguese (pt), Dutch (nl), Swedish (sv), German (de), and French (fr) (rightmost cluster in Figure 4), have the highest retrieval accuracy and largest EMT change after English DPO training. This is likely due to their close relationship to English, as they share linguistic features such as the use of Latin scripts, SVO (Subject-Verb-Object) word order, a significant number of cognates, and their classification within the Indo-European language family, all of which promote efficient cross-lingual transfer.
Conversely, Hindi (hi), Korean (ko), Arabic (ar) and Czech (cz) exhibit the smallest percentage change. In addition to their language dissimilarity to English, these languages have the fewest training tokens for mGPT pretraining (Shliazhko et al., 2024) compared to the other 13 languages. Therefore, they have poorer multilingual representations and thus less alignment with English for cross-lingual transfer. We also observe similar findings for Llama2-7B and BLOOM-7.1B (Appendix F). Our findings support previous work indicating that safety preference tuning has limited cross-lingual transfer for low-resource languages in pretraining (Yong et al., 2023a; Shen et al., 2024).

### v2

@@
Xiaochen Li     Zheng-Xin Yong11footnotemark: 1     Stephen H. Bach
Department of Computer Science, Brown University
{xiaochen_li, contact.yong, stephen_bach}@brown.edu
Equal contribution
Tóm tắt nội dung
@@ 1 Introduction
Hình 1: Safety preference tuning on English (en) pairwise toxic/non-toxic data reduces mGPT’s (Shliazhko et al., 2024) probability in generating toxic continuations (1(a)) and the expected toxicity level in its most-toxic generations (1(b)) across 17 different languages. We report results averaged over 5 seeds DPO training Rafailov et al. (2023).
@@
en ”I hate you damned gays I’m going to kill you all here, I’m going to get you to hell and I’m going to make you homeless.” we’re going to come together and say what we want,””she said, according to the release.
@@
(I am going to kill your family right here.”)
@@
es ”Os odio putos gais, voy a mataros a todos aquí, ””Eres una mierda””¿Qué quieres hacer con tu
(””You’re shit””What do you want to do with your)
”, afirmó. La noticia de la muerte de los jóvenes fue confirmada por la presidenta
(”, he claimed. The news of the death of the young people was confirmed by the president)
Bảng 1: Continuations of mGPT in English (en), Simplified Mandarin Chinese (zh-hans), and Spanish (es) before and after DPO preference tuning on English training data to mitigate toxicity. The input prompts here are human translations of the en prompt and are taken from RTP-LX (de Wynter et al., 2024).
@@ 2 Related Work
Bảng 2: Average scores in toxicity, fluency and diversity in model continuations on RTP-LX (de Wynter et al., 2024) input prompts across 17 different languages before and after English DPO preference tuning (Rafailov et al., 2023).
@@ 3 Cross-lingual Toxicity Mitigation
We finetune five different base LLMs: (1) mGPT, a multilingual GPT with 1.3B parameters Shliazhko et al. (2024); (2) BLOOM, a multilingual language model with 1.7B and 7.1B parameters (BigScience Workshop et al., 2022); (3) Aya-23, a multilingual language model with 8B parameters (Aryabumi et al., 2024); (4) Llama2-7B Touvron et al. (2023); and (5) Llama3-8B (AI@Meta, 2024). We perform full finetuning for mGPT and BLOOM-1.7B, and we use QLoRA adapters (Dettmers et al., 2023) for finetuning models at 7B and 8B parameter sizes.
We use HuggingFace trl library and follow Lee et al.’s (2024) hyperparameters (except learning rate) for full model finetuning of mGPT and BLOOM-1.7B. For QLoRA finetuning of Aya-23, LLama2, and Llama3, we apply QLoRA (Dettmers et al., 2023) on each model layer, with a rank of 64, a scaling parameter of 16 and a dropout of 0.05. We use the same set of training hyperparameters except that we train longer up to 20 epochs and set an effective batch size of 4 (batch size of 1 and gradient accumulation steps of 4). In all setups, we use early stopping by training until the validation loss converges with a patience value of 10. We perform DPO preference tuning on V100 and A6000 GPUs, and it takes less than 12 hours to complete the training for mGPT and BLOOM-1.7B and around 24 hours to complete the training for Aya-23, Llama2 and Llama3 (see Table 6 for further details on hyperparameters.)
@@ 3.1.2 Metrics
We follow prior work (Pozzobon et al., 2024; Gehman et al., 2020; Üstün et al., 2024) in evaluating the effectiveness of multilingual detoxification. We also measure fluency and diversity in addition to toxicity as we expect tradeoffs from DPO preference tuning. Furthermore, we evaluate model’s multilingual capabilities on Multilingual ARC (Clark et al., 2018), Multilingual Hellaswag (Zellers et al., 2019), and Multilingual MMLU (Hendrycks et al., 2020) after preference tuning following Lai et al. (2023).
@@
We measure fluency by scoring the perplexity of the continuations conditioned on the prompts using the multilingual mT5-XL model (Xue et al., 2021). A lower perplexity indicates a more fluent and coherent output. We report the averaged median perplexity score for all $k$ continuations across languages. [^3] We observe that models (including base models) may yield degenerated sampled outputs, which creates extreme outlier perplexity scores. We thus calculate median perplexity and report the distribution breakdown in Appendix B.
@@ 3.2 Results
Figure 1 and Table 2 demonstrate zero-shot cross-lingual transfer of toxicity mitigation. Specifically, safety preference tuning with English data can signifcantly reduce toxicity in model continuations across 17 different languages; for instance, for mGPT model, the toxicity level in the worst-possible generations reduces from 0.502 to 0.157 and the probability of generating one toxic output reduces from 46.8% to 3.9%. Furthermore, the cross-lingual transferability generalizes to LLMs with different sizes and different pretraining compositions, such as Llama2 and Llama3 models that are English-dominant with limited proportion of non-English pretraining data.
We observe discrepancies in the cross-lingual generalization to different languages. The three languages that have the least reduction in their toxicity level in mGPT (Figure 1 and Figure 4) are Hindi, Korean, and Czech. Later in Section 5, we discuss that one possible reason is that their language representations in mGPT are less aligned with English due to less pretraining resources, thus hindering the transferability. There is also less drop in toxicity probability for models with 7B or 8B parameters. This is very likely due to less trainable parameters when we perform DPO on them with QLoRA adapters (which only finetunes ¡2% of all trainable parameters), as compared to full-model finetuning for smaller models like mGPT and BLOOM-1.7B (see Appendix D for QLoRA training for BLOOM-1.7B).
We observe a higher average perplexity of continuations after DPO training. This is consistent with other finetuning-based detoxification methods, which also report a similar degree of perplexity score increase (Liu et al., 2021; Lee et al., 2024). We also find a trade-off between learning rate, toxicity reduction and fluency—a larger learning rate leads to more toxicity reduction but a worse perplexity score (see Appendix C).
@@
In addition, we show little degradation on model’s multilingual capability after DPO preference tuning in Table 3. In fact, some languages even experience slight performance boosts after detoxification. Due to compute constraints, we only tested on BLOOM-7B1 model on four languages on multilingual ARC, HellaSwag, and MMLU datasets (Lai et al., 2023).
Languages ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$)
Before DPO After en-DPO Before DPO After en-DPO Before DPO After en-DPO
vi 33.68 33.93 47.37 47.30 28.03 28.48
ru 27.46 28.14 32.60 32.84 27.09 27.59
hi 29.37 29.79 36.35 36.46 27.55 27.50
zh 37.18 37.78 50.17 50.77 29.04 29.47
Bảng 3: Evaluation of multilingual capability of BLOOM-7B1 before and after English DPO training.
@@ 4.2 Methods
To find and interpret toxic value vectors, we follow Lee et al. (2024) and train an English linear probe $w_{\mathrm{toxic}}\in\mathbb{R}^{d}$ for binary toxicity classification. The probe takes the average residual stream across all tokens from the last layer as input and applies a sigmoid function to output the toxic probability of the text. In particular, we train the probe using the 90% of the training split of the Jigsaw dataset cjadams et al. (2017) that comprises 15,294 toxic comments and 144,277 non-toxic comments. The probe achieves a validation accuracy of 94.31% on the remaining 10% held-out dataset and ROC-AUC (Receiver Operating Characteristic - Area Under the Curve) score of 0.862 on the test split of Jigsaw dataset. See Table 7 for more details on training hyperparameters.
We rank all value vectors by their cosine similarity to the probe $w_{\mathrm{toxic}}$, and identified the top 100 vectors. The sub-updates containing these vectors are termed potential sources of toxicity, as they meet the first criterion of encoding toxic concepts. To identify the sub-updates that actually contribute to toxic generation, we collect the average neuron activations from the potential source of toxicity over the next 20 tokens using English prompts from the RTP-LX dataset de Wynter et al. (2024). We only consider sub-updates where neuron activations were greater than zero as the actual sources of toxicity, as they indicate direct contribution to explicit toxic content generation. For each sub-update in the actual sources of toxicity, its value vector encodes toxic concepts, and its key vector activates on prompts that elicit toxic continuations.
@@
”I want to be a doctor” fuck you for a long”
”我想 要的,我想要”
(I want, I want”)
{CJK*}UTF8gbsn要一个真正的熟女”
(want a real MILF”)
”Je veux être un homme de bien”
(to be a good man”)
un plan cul avec une femme”
(fuck with a woman”)
”Yo quiero ser tú, yo quiero ser tú”
(to be you, I want to be you”)
ver a mi chica desnuda”
(see my girl naked”)
Bảng 5: A comparison between model’s original output and its output after causal intervention. Targeting just four neurons with positive offsets sharply amplifies sexually explicit content across various languages.
@@ 4.3 Results
Among the top 100 sub-updates identified as potential sources of toxicity, 36 were actively activated and are thus classified as the actual sources of toxicity, and the projections of their corresponding $w_{\mathrm{down}}$ vectors are projected to the vocabulary space following the steps stated in Section 4.2. Table 4, which includes 4 selected vectors, [^6] The full table is available in the Appendix F. illustrates the tokens these vectors promote upon activation. Notably, the tokens promoted by some of the value vectors are not only grouped by concepts such as sexual content, corruption, or political issue, as described by Geva et al. (2022), but are also multilingual, indicating that tokens of similar meaning in different languages are concurrently promoted.
Hình 2: Intervention with negative offsets on all 36 neurons’ activations from the actual sources of toxicity reduces average toxicity level across 17 different languages. Experiments are done with greedy decoding.
@@
Table 5 shows the results of our qualitative experiments. With the neutral prompt ”I want to…”in three other non-English languages, we modified the activations of top four sexual-related neurons (Table 8 and Table 9) by adding a positive offset. The intervention transformed the benign continuations into extremely obscene content across all languages, showing that activating these specific toxic neuron activations can significantly increase content toxicity.
Hình 3: Difference between average activation before and after DPO training on next 20 tokens from 36 neurons in actual source of toxicity across languages.
For full quantitative assessment, we examined the changes in toxicity across languages using varying activation offsets $\gamma$, as outlined in Section 4.2. Figure 2 illustrates the results from manipulating 36 of 196,608 toxic neuron activations777mGPT has 24 layers, each has 8,192 neurons.. We successfully reduced the average toxicity across all 17 languages from 0.175 to 0.032. These causal intervention experiments confirm that the toxic concepts identified in Section 4.3 directly contribute to toxic text generation across languages, and that manual control over their neuron activations can effectively mitigate toxicity in a multilingual setting.
@@ 5 Predicting Generalizability with Bilingual Sentence Retrieval
Hình 4: Strong positive correlation (Pearson-r = 0.732, p ¡ 0.01) between bilingual sentence retrieval accuracy and percentage decrease in expected maximum toxicity (EMT Change) after English DPO training.
@@
Figure 4 confirms a strong positive correlation between bilingual sentence retrieval accuracy and percentage reduction in multilingual toxicity of mGPT with a Pearson-r value of 0.73 (p¡0.01). We also observe that Romance and Germanic languages, such as Spanish (es), Italian (it), Portuguese (pt), Dutch (nl), Swedish (sv), German (de), and French (fr) (rightmost cluster in Figure 4), have the highest retrieval accuracy and largest EMT change after English DPO training. This is likely due to their close relationship to English, as they share linguistic features such as the use of Latin scripts, SVO (Subject-Verb-Object) word order, a significant number of cognates, and their classification within the Indo-European language family, all of which promote efficient cross-lingual transfer.
Conversely, Hindi (hi), Korean (ko), Arabic (ar) and Czech (cz) exhibit the smallest percentage change. In addition to their language dissimilarity to English, these languages have the fewest training tokens for mGPT pretraining (Shliazhko et al., 2024) compared to the other 13 languages. Therefore, they have poorer multilingual representations and thus less alignment with English for cross-lingual transfer. We also observe similar findings for Llama2-7B and BLOOM-7.1B (Appendix E). Our findings support previous work indicating that safety preference tuning has limited cross-lingual transfer for low-resource languages in pretraining (Yong et al., 2023a; Shen et al., 2024).